# Hybrid quantum–classical matrix-product state and Lanczos methods for electron–phonon systems with strong electronic correlations: Application to disordered systems coupled to Einstein phonons

### v1

@@ IV.2 Many electrons, $V=0$: Decay of CDW order
The evolution of the $\mathcal{O_{\rm{CDW}}}$ averaged over $4000$ trajectories is shown in Fig. 10. For $\gamma=0$, $\mathcal{O_{\rm{CDW}}}(t)$ saturates quickly and stays finite at all times, implying Anderson localization. By making the coupling to the phonons finite, though, localization is not stable anymore. The imbalance $\mathcal{O_{\mathrm{CDW}}}$ decays following a power law. We will discuss the exponents of the power-law decay together with the interacting case in Sec. V.
@@ V Stability of MBL
When we introduce interactions between electrons, when uncoupled to phonons, finite systems exhibit many-body localization (MBL) for large disorder strengths [dAlessio2016, Gogolin2016, Deutsch2018, Mori2018]. The stability of MBL in the presence of a bath has been studied in various scenarios both theoretically [Nandkishore2015, Sarang2017, Luitz2017, Katharine2017, Pollmann2020, Abanin12021] and experimentally [Bloch2019, Greiner2023]. In some cases, the bath delocalizes MBL [Sarang2017, Luitz2017, Pollmann2020, Abanin12021] (on accessible time scales and system sizes), whereas in other cases, MBL can even localize the bath [Nandkishore2015, Katharine2017] via a proximity effect. In this section we study the stability of finite-size MBL when coupled to a classical bath consisting of classical dispersion-less phonons.
@@ V.1 Decay of charge-density-wave states
We consider a large disorder strength $W=$8\text{\,}\mathrm{\mathit{t}_{0}}$\$ and similar to the previous case of $V=0$ from Sec. IV, we time evolve the CDW state (see Eq. 7) for different values of $\gamma$ [Abanin2023]. Here, the Lanczos-MTE is used for $L=14$ and the evolution of $\mathcal{O}_{\mathrm{CDW}}(t)$ is averaged over $4000$ trajectories.
@@ V.2 Power-law decay
As a significant feature, we observe that $\mathcal{O}_{\mathrm{CDW}}$ decays asymptotically like a power-law for all the combinations of system parameters $\gamma>0$. We observe, that adding interactions to the electron system leads to a faster decay of $\mathcal{O}_{\text{CDW}}$ at small $\gamma$ as compared to the $V=0$ case. At large $\gamma$ this effect is strongly reduced: In fact, we observe that the decay of $\mathcal{O}_{\mathrm{CDW}}$ is much reduced when $V>0$ for large $\gamma$. We will provide a qualitative explanation of these trends in the next section.
In the single-particle case, we clearly observe subdiffusion at finite $\gamma$ and $W>0$, with dynamical exponents $z\leq 0.5$ for large $W=$8\text{\,}\mathrm{\mathit{t}_{0}}$\$. While it is not obvious that this behavior carries over to the many-body case, there are nevertheless suggestions for a connection between the dynamical exponent $z$ and the change in the behavior of $\mathcal{O}_{\text{CDW}}$ (or more generally, density autocorrelations) [Luitz2017a]. This would suggest sub-diffusive dynamics even in the many-body case. Whether such behavior can persist up to asymptotically long times is an open question left for future research.
@@ V.3 Discussion and Interpretation
In this section we will provide a discussion of these qualitative observations. Generally, without the MTE approximation, a finite $\gamma$ introduces interactions between the electrons. One may therefore attempt to integrate out the phonons, mapping everything to an effective electronic model. This, however, will be insufficient to lead to the observed fast decay of $\mathcal{O}_{\text{CDW}}$ as an actual energy transfer to the phononic subsystem appears to be crucial for delocalization. The rate of that transfer is controlled by $1/\gamma$. Nonetheless, at weak $\gamma$, it is not surprising that larger electronic interactions accelerate the decay of $\mathcal{O}_{\text{CDW}}$ as this trend is already seen in the absence of a coupling to any phonons [Luitz2016a, Luitz2017a].
In the regime of large $\unit\gamma>\unit{}$, without resorting to the classical and mean-field approximations of MTE, increasing $\gamma$ will make the physics more local and in the extreme case produce small polarons. These quasiparticles have a smaller bandwidth than the bare electron and will therefore generally see a large effective disorder potential (for a more detailed discussion, see [Bronold2002, Bronold2004]). This suggests that at fixed disorder strength $W$, there are two competing trends: increasing $\gamma$ favors the energy transfer to the phonons and hence relaxation of density inhomogeneities but at the same time leads to less mobile particles and thus in itself to localization. Our numerical observations from Figures 10 and 11 are consistent with this picture.
There is another path towards understanding the delocalization due to the Holstein coupling to phonons that is particularly evident in the MTE approximation. The fluctuating phonon degrees of freedom couple directly to the electronic density $\langle\hat{n}_{\ell}\rangle$ and hence themselves induce disorder potential dynamically [DiSante2017]. On average, over time and trajectories, these average out, yet can cancel out the static disorder potential instantaneously.
@@
Figure 12: Sketches explaining relaxation processes induced by coupling to classical phonons in a single trajectory/realization. 12 Examplary setup of the chain for the CDW initial state. On each site $\ell$ there is a random disorder field $\epsilon_{\ell}$. On top of this static disorder the electron–phonon coupling imposes a dynamical quasi-disorder field, depending on the classical coordinates $x_{\ell}$. 12 Local relaxation process of an electron induced by the dynamical disorder fluctuations. As the energy difference between neighboring sites becomes small the electron delocalizes over both sites on a much shorter timescale than the fluctuations in the disorder, therefore resonant tunneling to to the next site is possible. 12 Sketch created from real data of a single realization for the parameters of Fig. 8. We choose a large value of $\gamma=$5\text{\,}\mathrm{\mathit{t}_{0}}$\$ to exaggerate the process that we want to highlight. On a selected lattice site $\ell=46$ and over a selected time frame, we show the site occupation $\langle\hat{n}_{\ell}\rangle$ and the on-site energy differences $\Delta E_{\ell-1,\ell}$ ($\Delta E_{\ell,\ell+1}$) induced by the time-dependent disorder between the left (right) lattice sites. As a guide to the eye, we have added circles at some of the points in time where $\langle\hat{n}_{\ell}\rangle$ jumps due to $\Delta E_{\ell-1,\ell}=0$ or $\Delta E_{\ell,\ell+1}=0$. The values of $\langle\hat{n}_{\ell}\rangle$ and $\Delta E_{\ell-1,\ell}$ ($\Delta E_{\ell,\ell+1}$) are rescaled for better visibility by setting the maximum value of each curve over the inspected interval to one.
This leads to a classical picture of resonant local relaxation processes. We consider a single trajectory in the limit of $\unit\gamma>\unit{}$. The total effective disorder at lattice site $\ell$ at time $t$ is $\epsilon_{\mathrm{total},\ell}=\epsilon_{\ell}-\sqrt{2}\gamma x_{\ell}(t)$. We observe that whenever $\epsilon_{\mathrm{total},\ell}$ on two neighboring sites is nearly equal, the electron can resonantly tunnel into a previously empty site. A sketch is shown in Fig. 12. Since the electron sector undergoes much faster dynamics than the disorder landscape at small $\omega_{0}$, changes in the potential landscape due to the classical phonons may enable local relaxation processes. The presence of repulsive interactions detunes these direct transition, hence at large $\gamma$, increasing $V$ leads to slower relaxation.
@@ VI Conclusion
There are several interesting extensions and open questions to our work. On the technical side, the application of TEBD-MTE to dynamical quantities — e.g., spectral functions, and conductivities — seems promising. Moreover, TEBD can be replaced by other time-propagating schemes for matrix-product states [Paeckel2019]. In the long run, a combination of matrix-product state methods and coupled-trajectory methods (such as, for instance multiconfigurational Ehrenfest [Shalashilin2009, Shalashilin2010]) may be lead to much more reliable results than MTE.
As for the physics of interacting electrons in the presence of disorder, few numerical studies have been conducted that include phonons as a generic relaxation channel in a solid-state environment while treating electronic correlations at a finite filling adequately [Bronold2002, Bronold2004, Berciu2010]. Many interesting questions present themselves, such as a full characterization of transport [Miladic2025], establishing a connection to variable-range hopping, and the extension to acoustic phonons, to name a few.
Simulation codes and the data shown in the figures is publicly available on Zenodo [this_zenodo].
Note added: During the final stages of preparing this manuscript, we became aware of a study of CDW dynamics in Holstein chains using a classical approximation to the phonon sector [Jang2025].
@@ Appendix A Behavior of single trajectories in MTE
Single trajectories under MTE dynamics may undergo chaotic dynamics. It turns out that single trajectories show a sensitivity to initial conditions as one would expect from a classically chaotic system constituent of only classical degrees of freedom. While the Ehrenfest dynamics make up a mixed quantum–classical system, the quantum subsystem acts like a time-dependent driving of the harmonic oscillators in the classical part of the system. Indeed, driving a classical system time-dependently is one of the simplest ways to create sensitivity to initial conditions [Schuster1995].
This is reflected in the dependence of trajectories on initial conditions and numerical control parameters. As Figures 3 and 4 shows, the electronic energy develops a large deviation between originally nearby trajectories, independent of the initial deviation in the coordinate at a single site. Moreover, we see that the initial state dependence of the classical system also leads to large deviations in the electronic observables, see Fig. 13 where we have investigated the difference of the electronic energy $\delta E_{\mathrm{el}}=|E_{\mathrm{el}}(0)-E_{\mathrm{el}}(\delta x)|$ between slightly perturbed trajectories. Here, $\delta x$ is a small perturbation acting on the harmonic oscillator position coordinate $x_{\ell}$ at site $\ell=50$ and time $t=0$ like so, $x_{\ell}(\delta x)=x_{\ell}(0)+\delta x$. We observe that already a small perturbation $\delta x=10^{-8}$ is enough to have large effects on the dynamics of the system, as would be expected in a classically chaotic system.
@@ Appendix B Convergence of TEBD-MTE for large systems
Figure 13: Dynamics in a single trajectory: Decay of the CDW in the non-interacting system (L=100{L=100}, V=0{V=0}, $\gamma=$0.5\text{\,}\mathrm{\mathit{t}_{0}}$\$}, $W=$8\text{\,}\mathrm{\mathit{t}_{0}}$\$}, $\omega_{0}=$0.1\text{\,}\mathrm{\mathit{t}_{0}}$\$}). We compare trajectories with slightly different initial conditions: The trajectories only vary in their initial condition at site $\ell=50$ by $x_{\ell}(\delta x)=x_{\ell}(0)+\delta x$. The data sets show the resulting difference in the electronic energy $\delta E_{\mathrm{el}}=|E_{\mathrm{el}}(0)-E_{\mathrm{el}}(\delta x)|$. The large increase of $\delta E_{\mathrm{el}}$ is seen over a wide range of investigated parameter regimes.
[…]
Figure 14: Dependence of the TEBD-MTE method on the discarded weight $\delta$ and the time step $\Delta t$ in an interacting system ($L=28$, $V=$2\text{\,}\mathrm{\mathit{t}_{0}}$\$, $\omega_{0}=$0.1\text{\,}\mathrm{\mathit{t}_{0}}$\$, $\gamma=$0.4\text{\,}\mathrm{\mathit{t}_{0}}$\$, $W=0$). We plot the deviation $\delta\mathcal{O}_{\mathrm{CDW}}$ when 14 varying $\delta$ at a fixed time step $\Delta t=$0.01\text{\,}{\mathrm{\mathit{t}_{0}}}^{-1}$\$, comparing with a reference set using $\delta=10^{-10}$ and 14 varying $\Delta t$ at fixed cutoff $\delta=10^{-10}$, comparing with reference set using $\Delta t=$0.01\text{\,}{\mathrm{\mathit{t}_{0}}}^{-1}$\$. The results are averaged over $N_{\mathrm{traj}}=4000$ trajectories.

### v2

@@ IV.2 Many electrons, $V=0$: Decay of CDW order
The evolution of $\mathcal{O_{\rm{CDW}}}$ averaged over $4000$ trajectories is shown in Fig. 10. For $\gamma=0$, $\mathcal{O_{\rm{CDW}}}(t)$ saturates quickly and stays finite at all times, implying Anderson localization. By making the coupling to the phonons finite, though, localization is not stable anymore. The imbalance $\mathcal{O_{\mathrm{CDW}}}$ decays following a power law. We will discuss the exponents of the power-law decay together with the interacting case in Sec. V.
@@ V Stability of MBL
When we introduce interactions between electrons, when uncoupled to phonons, finite systems exhibit many-body localization (MBL) for large disorder strengths [16, 30, 19, 69]. The stability of MBL in the presence of a bath has been studied in various scenarios both theoretically [71, 32, 59, 40, 107, 54] and experimentally [84, 55]. In some cases, the bath delocalizes MBL [32, 59, 107, 54] (on accessible time scales and system sizes), whereas in other cases, MBL can even localize the bath [71, 40] via a proximity effect. In this section we study the stability of finite-size MBL when coupled to a classical bath consisting of classical dispersion-less phonons.
@@ V.1 Decay of charge-density-wave states
We consider a large disorder strength $W=$8\text{\,}\mathrm{\mathit{t}_{0}}$\$ and similar to the previous case of $V=0$ from Sec. IV, we time evolve the CDW state (see Eq. 7) for different values of $\gamma$ [8]. Here, the Lanczos-MTE is used for $L=14$ and the evolution of $\mathcal{O}_{\mathrm{CDW}}(t)$ is averaged over $4000$ trajectories.
@@ V.2 Power-law decay
As a significant feature, we observe that $\mathcal{O}_{\mathrm{CDW}}$ decays asymptotically with a power law for all values of $\gamma$ considered. We observe, that adding interactions to the electron system leads to a faster decay of $\mathcal{O}_{\text{CDW}}$ at small $\gamma$ as compared to the $V=0$ case. At large $\gamma$ this effect is strongly reduced: In fact, we observe that the decay of $\mathcal{O}_{\mathrm{CDW}}$ is much reduced when $V>0$ for large $\gamma$. We will provide a qualitative explanation of these trends in the next section.
In the single-particle case, we clearly observe subdiffusion at finite $\gamma$ and $W>0$, with dynamical exponents $z\leq 0.5$ for large $W=$8\text{\,}\mathrm{\mathit{t}_{0}}$\$. While it is not obvious that this behavior carries over to the many-body case, there are nevertheless suggestions for a connection between the dynamical exponent $z$ and the change in the behavior of $\mathcal{O}_{\text{CDW}}$ (or more generally, density autocorrelations) [61]. This would suggest sub-diffusive dynamics even in the many-body case. Whether such behavior can persist up to asymptotically long times is an open question left for future research.
@@ V.3 Discussion and Interpretation
In this section we will provide a discussion of these qualitative observations. Generally, without the MTE approximation, a finite $\gamma$ introduces interactions between the electrons. One may therefore attempt to integrate out the phonons, mapping everything to an effective electronic model. This, however, will be insufficient to lead to the observed fast decay of $\mathcal{O}_{\text{CDW}}$ as an actual energy transfer to the phononic subsystem appears to be crucial for delocalization. The rate of that transfer is controlled by $1/\gamma$. Nonetheless, at weak $\gamma$, it is not surprising that larger electronic interactions accelerate the decay of $\mathcal{O}_{\text{CDW}}$ as this trend is already seen in the absence of a coupling to any phonons [60, 61].
In the regime of large $\unit\gamma>\unit{}$, without resorting to the classical and mean-field approximations of MTE, increasing $\gamma$ will make the physics more local and in the extreme case produce small polarons. These quasiparticles have a smaller bandwidth than the bare electron and will therefore generally see a large effective disorder potential (for a more detailed discussion, see [13, 12]). This suggests that at fixed disorder strength $W$, there are two competing trends: increasing $\gamma$ favors the energy transfer to the phonons and hence relaxation of density inhomogeneities but at the same time leads to less mobile particles and thus in itself to localization. Our numerical observations from Figures 10 and 11 are consistent with this picture.
There is another path towards understanding the delocalization due to the Holstein coupling to phonons that is particularly evident in the MTE approximation. The fluctuating phonon degrees of freedom couple directly to the electronic density $\langle\hat{n}_{\ell}\rangle$ and hence themselves induce disorder potentials dynamically [20]. On average, over time and trajectories, these average out, yet can cancel out the static disorder potential instantaneously.
@@
Figure 12: Sketches explaining relaxation processes induced by coupling to classical phonons in a single trajectory/realization. 12 Examplary setup of the chain for the CDW initial state. On each site $\ell$ there is a random disorder field $\epsilon_{\ell}$. On top of this static disorder the electron–phonon coupling imposes a dynamical quasi-disorder field, depending on the classical coordinates $x_{\ell}$. 12 Local relaxation process of an electron induced by the dynamical disorder fluctuations. As the energy difference between neighboring sites becomes small the electron delocalizes over both sites on a much shorter timescale than the fluctuations in the disorder, therefore resonant tunneling to the next site is possible. 12 Sketch created from real data of a single realization for the parameters of Fig. 8. We choose a large value of $\gamma=$5\text{\,}\mathrm{\mathit{t}_{0}}$\$ to exaggerate the process that we want to highlight. On a selected lattice site $\ell=46$ and over a selected time frame, we show the site occupation $\langle\hat{n}_{\ell}\rangle$ and the on-site energy differences $\Delta E_{\ell-1,\ell}$ ($\Delta E_{\ell,\ell+1}$) induced by the time-dependent disorder between the left (right) lattice sites. As a guide to the eye, we have added circles at some of the points in time where $\langle\hat{n}_{\ell}\rangle$ jumps due to $\Delta E_{\ell-1,\ell}=0$ or $\Delta E_{\ell,\ell+1}=0$. The values of $\langle\hat{n}_{\ell}\rangle$ and $\Delta E_{\ell-1,\ell}$ ($\Delta E_{\ell,\ell+1}$) are rescaled for better visibility by setting the maximum value of each curve over the inspected interval to one.
This leads to a classical picture of resonant local relaxation processes. We consider a single trajectory in the limit of $\unit\gamma>\unit{}$. The total effective disorder at lattice site $\ell$ at time $t$ is $\epsilon_{\mathrm{total},\ell}=\epsilon_{\ell}-\sqrt{2}\gamma x_{\ell}(t)$. We observe that whenever $\epsilon_{\mathrm{total},\ell}$ on two neighboring sites is nearly equal, the electron can resonantly tunnel into a previously empty site. A sketch is shown in Fig. 12. Since the electronic sector undergoes much faster dynamics than the disorder landscape at small $\omega_{0}$, changes in the potential landscape due to the classical phonons may enable local relaxation processes. The presence of repulsive interactions detunes these direct transition, hence at large $\gamma$, increasing $V$ leads to slower relaxation. This is illustrated in Fig. 13.
[…]
@@ VI Conclusion
There are several interesting extensions and open questions to our work. On the technical side, the application of TEBD-MTE to dynamical quantities — e.g., spectral functions, and conductivities — seems promising. Moreover, TEBD can be replaced by other time-propagating schemes for matrix-product states [75]. In the long run, a combination of matrix-product state methods and coupled-trajectory methods (such as, for instance multiconfigurational Ehrenfest [89, 90]) may be lead to much more reliable results than MTE.
As for the physics of interacting electrons in the presence of disorder, few numerical studies have been conducted that include phonons as a generic relaxation channel in a solid-state environment while treating electronic correlations at a finite filling adequately [13, 12, 5]. Many interesting questions present themselves, such as a full characterization of transport [66], establishing a connection to variable-range hopping, and the extension to acoustic phonons, to name a few.
Simulation codes and the data shown in the figures is publicly available on Zenodo [65].
Note added: During the final stages of preparing this manuscript, we became aware of a study of CDW dynamics in Holstein chains using a classical approximation to the phonon sector [43].
@@ Appendix A Behavior of single trajectories in MTE
Figure 14: Dynamics in a single trajectory: Decay of the CDW in the non-interacting system (L=100{L=100}, V=0{V=0}, $\gamma=$0.5\text{\,}\mathrm{\mathit{t}_{0}}$\$}, $W=$8\text{\,}\mathrm{\mathit{t}_{0}}$\$}, $\omega_{0}=$0.1\text{\,}\mathrm{\mathit{t}_{0}}$\$}). We compare trajectories with slightly different initial conditions: The trajectories only vary in their initial condition at site $\ell=50$ by $x_{\ell}(t=0)=x_{\ell}(t=0)+\delta x$. The data sets show the resulting difference in the electronic energy $\delta E_{\mathrm{el}}=|E_{\mathrm{el}}(0)-E_{\mathrm{el}}(\delta x)|$. The large increase of $\delta E_{\mathrm{el}}$ is seen over a wide range of investigated parameter regimes.
Single trajectories under MTE dynamics may undergo chaotic dynamics. It turns out that single trajectories show a sensitivity to initial conditions as one would expect from a classically chaotic system constituent of only classical degrees of freedom. While the Ehrenfest dynamics make up a mixed quantum–classical system, the quantum subsystem acts like a time-dependent driving of the harmonic oscillators in the classical part of the system. Indeed, driving a classical system time-dependently is one of the simplest ways to create sensitivity to initial conditions [88].
This is reflected in the dependence of trajectories on initial conditions and numerical control parameters. As Figures 3 and 4 shows, the electronic energy develops a large deviation between originally nearby trajectories, independent of the initial deviation in the coordinate at a single site. Moreover, we see that the initial state dependence of the classical system also leads to large deviations in the electronic observables, see Fig. 14 where we have investigated the difference of the electronic energy $\delta E_{\mathrm{el}}=|E_{\mathrm{el}}(0)-E_{\mathrm{el}}(\delta x)|$ between slightly perturbed trajectories. Here, $\delta x$ is a small perturbation acting on the harmonic oscillator position $x_{\ell}$ at site $\ell=50$ and time $t=0$ , changing the initial condition to $x_{\ell}(t=0)=x_{\ell}(t=0)+\delta x$. We observe that already a small perturbation $\delta x=10^{-8}$ is enough to have large effects on the dynamics of the system, as would be expected in a classically chaotic system.
@@ Appendix B Convergence of TEBD-MTE for large systems
[…]
Figure 15: Dependence of the TEBD-MTE method on the discarded weight $\delta$ and the time step $\Delta t$ in an interacting system ($L=28$, $V=$2\text{\,}\mathrm{\mathit{t}_{0}}$\$, $\omega_{0}=$0.1\text{\,}\mathrm{\mathit{t}_{0}}$\$, $\gamma=$0.4\text{\,}\mathrm{\mathit{t}_{0}}$\$, $W=0$). We plot the deviation $\delta\mathcal{O}_{\mathrm{CDW}}$ when 15 varying $\delta$ at a fixed time step $\Delta t=$0.01\text{\,}{\mathrm{\mathit{t}_{0}}}^{-1}$\$, comparing with a reference set using $\delta=10^{-10}$ and 15 varying $\Delta t$ at fixed cutoff $\delta=10^{-10}$, comparing with reference set using $\Delta t=$0.01\text{\,}{\mathrm{\mathit{t}_{0}}}^{-1}$\$. The results are averaged over $N_{\mathrm{traj}}=4000$ trajectories.